# VARX Granger Analysis: Modeling, Inference, and Applications

### v1

@@ 2.3 Debiased Deviance for L2 Regularization
where we selected ${\bf\Gamma}={\rm diag}({\bf R}_{xx})$ so that all variables are regularized equally regardless of their scale. The choice of $\gamma$ is discussed in the results section. This regularization introduces (purposefully) a bias in the estimate, and the deviance estimate has to be corrected to account for this bias [21]. The term that corrects the log-likelihood in case of L2 regularization is (derived in the Appendix):
@@
2.4 Effect size
Note that deviance increases linearly with $T$, that is, the statistical evidence increases with the length of the signals. The actual benefit of adding a link to the model is given by the log-ratio of the residual variances, and that remains constant with increasing signal duration. This log-ratio can be seen as a measure of effect size. Regularization can make this log-ratio negative and this is what the bias terms correct for. In this work, we take $\mathscrbf{D}^{de-biased}/T$ as a metric of effect size (e.g. in Fig. 9). A traditional definition of effect size in the context of reduced and full linear models is the coefficient of determination, or generalized R-square [22]:
$R^{2}:=\mathbf{1}-\exp\left(-\mathscrbf{D}^{de-biased}/T\right),$ (19)
where the exponential of a vector is interpreted element-wise.
@@ 2.5 Basis functions for the moving average filters
The filter length (number of parameters) used in AR filters is typically kept relatively short, to avoid over-fitting, reduce the odds of instability in the recursion, and because even a single delay can already represent an infinite impulse response. This is not the case for MA filters, where longer responses have to be modeled explicitly, resulting in a relatively large number of parameters, with a risk of over-fitting. We have found empirically that the corrections we introduced in the Deviance estimate for short signals (Eq. 18) do not work well when the filter lengths for ${\bf A}$ and ${\bf B}$ are very different. A solution to both these problems (imbalance in number of parameters and filter length) is to use basis functions for the ${\bf B}$ filters, following an approach used previously for TRFs [23, 24, 25]. In this formalism, we have:
@@
3 Results
Details on all results provided next can be found in the accompanying code repository https://github.com/lcparra/varx.
3.1 Test of model estimation on known model
To validate the estimation algorithm and code, we simulated a simple VARX model with two outputs and one input ($d_{y}=2,d_{x}=1$). The algorithm correctly recovers the AR and MA parameters (Fig.2). VARX model estimation is available as part of the econometric toolbox in Matlab, but is limited to instantaneous input $n_{b}=1$, i.e. no filtering of the input. When limiting the simulation to this case, the algorithms obtain similar results. Small variations are expected based on how the initial conditions (boundaries) are handled and numerical differences. In our implementation we omit from the estimate all samples that do not have a valid history. The code handles missing values (NaN) in the same way.
Figure 2: Comparison of estimated parameters to true parameters in a simple toy example. Here $x_{1}\rightarrow y_{2}$ and $y_{1}\rightarrow y_{2}$ were zero in the simulated data. In this simulation signals are generated with i.i.d innovation. We used $T=1000,\lambda=0$ and $n_{a}=3,n_{b}=1$ in both simulation and estimation.
3.2 Validation of p-values
To determine if the estimation computes correct $p$-values we simulated and VARX model with some channels set to zero, and others with non-zero values. We did this with a small and large simulated dataset, generated with normal i.i.d. innovations. We repeat the simulation 1000 times and determine how many times the zero channels report a $p<0.05$. We find a false discovery rate of approximately 0.05 for that specific null-channel, suggesting that $p$-values are correctly estimated (Fig. 3).
Figure 3: Validation of $p$-values in both a small and a larger model ($d_{y}=6$ and $d_{y}=60$). Significance is set of p<0.05, so we expect a false discovery rate of 0.05. Simulation here used with $d_{x}=1,n_{a}=2,n_{b}=2,T=1000,\lambda=0.0$. Filter coefficients for ${\bf B}$ were selected at random from unit variance normal, and ${\bf A}$ values were set to be ±$\pm 0.05$ with sign selected at random (this insured stable recursion in practice). Only two links are set to zero $A(:,2,2)=0,B(:,5,1)=0$, for these two, the $p$-value is correctly estimated at approximately $0.05$ (green).
3.3 Validation of p-values with L2 regularization
3.4 Example of using basis functions
@@
3.5 Example: Intracranial recordings in human
In our view, one of the main advantages of a VARX model is the factorization of the overall system response into an AR and MA portion, which separates the effect of the endogenous variables into an initial response that then reverberates in the recurrence of the dynamical system. To demonstrate this we analyze an example of an intracranial EEG recording in a 49-year-old male patient, included in [26]. The patient watched a total of 43.6 minutes of movie clips while recording iEEG and eyetracking. We focus on a subset of 50 electrodes covering visual brain areas, including the occipital cortex, fusiform face area and parahippocampal cortex. The neural data was preprocessed as described in [26] before extracting high-frequency broadband power (70-150Hz and downsampled to 60Hz). The exogenous "input" is a pulse train indicating the start of fixations, i.e. the moments when the subject receives new visual input. What is most evident is that the ${\bf B}$ response is shorter than the full impulse response ${\bf H}$. When estimating it directly or as ${\bf H}=(1-{\bf A})^{-1}{\bf B}$ we see that the two are very similar. This suggests that the VARX model decomposes the total response into a fast response followed by a prolonged response due to the recurrence in the brain network. The VARX model appears to be therefore a valid factorization of the full system response estimated traditionally.
Figure 6: Example of intracranial recording in humans. Here a VARX model is fit to Broad-band high-frequency activity during free viewing of 7 videos recorded in ($d_{y}=50$) electrodes. A total of 43.6 minutes of data were used at a sampling rate of 60Hz ($T=156,955$) from a single patient. (A) Effect size $R$ for the recurrent connectivity ${\bf A}$ between recording electrodes – in the language of neuroscience, this could be called "functional connectivity". (B) Effect size $R$ of fixation onset as exogenous variable on different electrodes. (C) ${\bf A}$ filter coefficients ($n_{a}=4$). Diagonal terms dominate. (D) ${\bf B}$ filter coefficients $n_{b}=30,\underline{n}=20$. $\lambda=0$
(E) System response estimated directly – in the language of neuroscience, this is the multivariate "temporal response function" (mTRF). (F) System response resulting from the VARX model estimate (Eq. 5). Data from [26].
3.6 Example: Physiological signals in human
Another dynamical system is the body, with multiple interconnected signals such as respiration, heart, pupil, and brain. Here we are analyzing data from [27] which showed bidirectional effects for several of these signals during various controlled conditions. We estimated a VARX model for this data suggesting possible directional effects between various physiological variables (Fig. 7). Pupil size and heart rate were measured in the experiment as they are conventional metrics of physiological arousal. As variables are added to the VARX model the connectivity structure is typically preserved. In this specific example using the controlled breathing task, we initially observe a bidirectional link between pupil size and heart rate, however this disappears once respiration is taken into account. Instead, it is explained by an effect of respiration on pupil size, together with the well-established bidirectional link between respiration and heart rate that is recovered in this data. The effect of saccades on pupil identified here is also well established. In general, adding variables can remove links – if the addition is a common cause, or add links – if the addition is a "collider". This is well established for i.i.d samples [28], and is no different for temporally correlated time sequence data.
Figure 7: Example of physiological signals in human. This data was collected while people carried out a controlled breathing intervention. In this case there was no exogenous stimulus, so we only fit a VAR model. Links in ${\bf A}$ are show if $p<0.001$. Here we had 26 minutes of data from compiled across multiple subjects samples at 25Hz ($T=26*60*25$). Data from [27].
3.7 Example: Sociology
As a final example, we present an analysis from the area of sociology. Here we analyze the history of union participation, and how it relates to strikes (Fig. 8A). The hypothesis we want to test is if strikes increase union membership in subsequent years. We assume that the unemployment rate can not be influenced by any of the union variables. It therefore is modeled as an exogenous input variable, whereas the number of workers in unions and on strike, as well as the number of strikes, can in principle all affect each other. VARX Granger analysis suggests (Fig. 8B) suggests that unemployment affects unionization, which in turn affects number of strikes, which obviously affects the number of workers on strike. These results depended on the choice of parameters $n_{a},n_{b},\lambda$. Only the effect of NumberOfStrikes $\rightarrow$ WorkersOnStrike was robust to parameter choice. What did not robustly emerge from this data is evidence for the initial hypothesis that strikes lead to an increase in union participation.
Figure 8: Example on union participation and strikes: (A) Historical data from the US. We treated the unemployment rate as an exogenous input in the VARX model, and the others as endogenous variables. (B) Significant effects in ${\bf A}$ and ${\bf B}$ are indicated in blue and red, respectively ($p<0.05$, $n_{a}=n_{b}=3,T=91,\lambda=0$). Note missing data around 1980, which was omitted during the estimation, including a 3-year history.
@@ 3.8 The case of a missing and superfluous variable
Finally, we wanted to evaluate the case where the model does not match the data generation process. We simulated three possible data generation processes with one input and two outputs ($d_{x}=1,d_{y}=2$). We simulated three cases, where the exogenous input $x$ (conditionaing variable) will be either a common cause, a collider, or an independent variable (see Figure 9). In all cases, the simulation implements a one-directional effect $y_{1}\rightarrow y_{2}$. We then measure how frequently we find $p<0.05$ for this path, i.e. the power of the test to identify a correct path, and how frequently we find $p<0.05$ for $y_{1}\leftarrow y_{2}$, i.e. the rate of false discovery. Note that conditioning on a collider is known to introduce spurious correlations [28]. We test this for data generated with both the equation (VARX) and output error models. The results in Figure 9 indicate the false discovery rate is correctly estimated at 0.05 in most scenarios, i.e. we are not finding causal effects above chance where there were none. This result holds regardless of whether $x$ was included as input (i.e. as a control variable with instant or delayed effect) or whether it did or did not have a true effect on $y_{1}$ and $y_{2}$ (common cause vs independent). Only when incorrectly modeling a collider as input, did we obtain spurious effects. Statistical power was improved when including the input to the model. In summary, there is no risk of false discovery when including input variables, even if they don’t have a true effect, except if they are actually affected by the internal variables $y$.
Despite using the VARX model, when the data was generated with an output error model, the false discovery remains limited at the target of 0.05. However, a common unobserved cause can generate spurious effects $y1\leftarrow y_{2}$. It has been suggested that running a Granger causality model on time-reversed data provides a control for this situation [29, 30]. We have found that running the model on time-reversed data results in spurious effects in all conditions tested here, so it is not clear to us how this can provide a remedy, and the caveat of an unobserved common cause remains when we are not directly observing endogenous variables, but only a noisy version of the internal dynamic.
@@
4 Discussion
Equation error versus output error models
The estimation of model parameters for a VARX model has a closed-form solution, thus being much faster than finding parameters for an output error model, which requires iterative algorithms. The gain in computational efficiency results from the assumption that $y(t)$ is observable. This may not be a good assumption in the case of brain signals measured across the skull, such as EEG/MEG where only a linear mixture, possibly with added noise, is observed. In that case, iterative algorithms are needed, but the Granger formalism can still be used with some effort [20]. In the VARX model, however, we do not need to assume that all internal activity is directly observable. Any unobserved activity is captured as innovation $e(t)$. We only need to be aware that any recurrent connectivity may be due to those unobserved common "causes". In particular symmetric effect sizes $R$ will be suggestive of such a missing variable. The role of the error is quite different in the two models. In the VARX (equation error) the error is an internal source of innovation driving the recurrent dynamic similar to the drive that comes from the input. The internal states are fully observable. In the output error model, the input entirely drives the system, and the error only affects the observations and is not injected into the dynamic.
Sensitivity to parameters
A caveat to all results above is that individual links can be sensitive to the model assumptions, namely, which variables are selected as endogenous (and can be affected by all others), and which variables are selected as exogenous (and cannot be affected). An example of that was the choice of the unemployment rate as exogenous to the dynamic of unions. The results can also depend on which endogenous variables are included, as we saw in the example of physiological signals. Results can also depend on the number of parameters $n_{a}$ and $n_{b}$ and regularization factor $\lambda$ (we saw this in the Union example). Although we did demonstrate this here, these parameters could be established with cross-validation.
Causality
In Granger’s original formalism [1], the error of the full and reduced model refers to one-dimensional signals where $y(t-1)$ is used in both cases and $x(t)$ is either used or omitted. If the error is significantly reduced by including $x(t)$ in the model, Granger argues that $x$causes" $y$. This interpretation is inaccurate for several reasons [31]. As we saw, when common causes are not observed (either as external input or internal variables) they can generate spurious links [32]. Similarly, including colliders can cause spurious links. All this is well explained by Pearl’s approach to causal inference [28]. Therefore one should not think of the Granger formalism as serious evidence for a causal graph without a well-justified prior graphical model [33]. In particular for large dimensional datasets such as brain data, where we only observe a tiny fraction of all the variables, the risk of unobserved common causes is much too large to take the resulting graph seriously as a causal graph. Nevertheless, asymmetries in the ${\bf A}$ matrix can be seen as evidence of temporal precedence suggestive of an asymmetric "information flow".
Non-stationarity
Toolboxes
There are several software tools to estimate VARX models. We are not aware of one that provides the Granger-Geweke test for significance. The MVGC toolbox [10] does not report results on exogenous variables, https://github.com/SacklerCentre/MVGC1. Implementations of VARX models in Matlab Econometrics Toolbox and in SAS for instance, make significance statements for individual delays but do not allow for the exogenous variable, i.e. $nb=1$. We welcome feedback if there other tools that that should be mentioned here. Our code emphasized computational efficiency to handle large datasets, with comparatively long ${\bf B}$ filters.
Stability
A word about ${\bf A}$ is in order. The AR filter $1/(1-{\bf A})$ can be unstable. We have not implemented any mechanism for this vectorial AR filter to remain stable. Lack of stability only manifests in the systems when computing the overall systems response $\mathcal{\bf H}$, which is not necessary during estimation for the ${\bf B},{\bf A}$ nor the calculation of statistical significance of each path (contrary to the output error model, where the recurrence has to be run back in time to estimated gradients, risking issues of stability (is this true?). We rarely encountered unstable AR estimates, and where we did, L2 regularization addressed the issues. But again, there is nothing in our formalism to ensure the stability of $\mathcal{\bf H}$.
1/f spectrum
5 Conclusion
The "multivariate temporal response function" (TRF) approach for modeling the response of dynamical systems to exogenous variables essentially boils down to a vectorial MA model. And, unlike the VAR model, TRF is not commonly examined in the Granger formalism. Although Geweke introduced exogenous variables into the Granger formalism [8], the clear connection to the VARX model is not yet widely recognized. We hope to have bridged this gap. Different models vary in their assumptions about how to break down the system’s overall response. When estimating with the VARX model, the data itself dictates how to factor the overall response into MA versus AR dynamics.
6 Acknowledgement
We would like to thank Vera E. Parra <veraeparra@gmail.com> for posing the question on union participation and compiling the corresponding datasets.
7 Appendix: Derivation of the De-biased Deviance
@@
Under the null hypothesis, the de-biased deviance difference can be shown to be asymptotically, as $T\rightarrow\infty$, chi-square distributed with $n$ degrees of freedom, where $n$ is the number of parameters removed in the reduced model and $T$ is the number of observations (i.e., column dimension of $\mathbf{Y}$) [18]. This asymptotic property requires that ridge estimator is consistent. To guarantee this, $\gamma$ must grow slower than $\mathcal{O}(T)$. In practice, fixing $\gamma$ using cross-validation typically guarantees this.

### v2

@@ 7.1 Debiased Deviance for L2 Regularization
where we selected ${\bf\Gamma}={\rm diag}({\bf R}_{xx})$ so that all variables are regularized equally regardless of their scale. The choice of $\gamma$ is discussed in the results section. This regularization introduces (purposefully) a bias in the estimate, and the deviance estimate has to be corrected to account for this bias [28]. The term that corrects the log-likelihood in case of L2 regularization is (derived in Section 7.6):
@@
7.2 Validation of p-values with L2 regularization
@@ 7.3 Basis functions for the moving average filters
The filter length (number of parameters) used in AR filters is typically kept relatively short, to avoid over-fitting, reduce the odds of instability in the recursion, and because even a single delay can already represent an infinite impulse response. This is not the case for MA filters, where longer responses have to be modeled explicitly, which can result in a relatively large number of parameters, with a risk of over-fitting. We have found empirically that the corrections we introduced in the Deviance estimate for short signals (Eq. 18) do not work well when the filter lengths for ${\bf A}$ and ${\bf B}$ are very different. A solution to both these problems (imbalance in number of parameters and filter length) is to use basis functions for the ${\bf B}$ filters, following an approach used previously for TRFs [29, 30, 31]. In this formalism, we have:
@@ 7.4 The case of a missing and superfluous variable
Here we want to evaluate the case where the model does not match the data generation process. We simulated three possible data generation processes with one input and two outputs ($d_{x}=1,d_{y}=2$). We simulated three cases, where the exogenous input $x$ (conditionaing variable) will be either a common cause, a collider, or an independent variable (see Figure 10). In all cases, the simulation implements a one-directional effect $y_{1}\rightarrow y_{2}$. We then measure how frequently we find $p<0.05$ for this path, i.e. the power of the test to identify a correct path, and how frequently we find $p<0.05$ for $y_{1}\leftarrow y_{2}$, i.e. the rate of false discovery. Note that conditioning on a collider is known to introduce spurious correlations [16]. We test this for data generated with both the equation (VARX) and output error models. The results in Figure 10 indicate the false discovery rate is correctly estimated at 0.05 in most scenarios, i.e. we are not finding causal effects above chance where there were none. This result holds regardless of whether $x$ was included as input (i.e. as a control variable with instant or delayed effect) or whether it did or did not have a true effect on $y_{1}$ and $y_{2}$ (common cause vs independent). Only when incorrectly modeling a collider as input, did we obtain spurious effects. Statistical power was improved when including the input to the model. In summary, there is no risk of false discovery when including input variables, even if they don’t have a true effect, except if they are actually affected by the internal variables $y$.
@@
7.5 Equation error versus output error model
An additional source of mismatch between the model and the data generation process is where error or the source of innovation originates. The VARX model in Eq. (1) is also called “equation error” model [1] because the error breaks the equality of the MA and AR terms. The equation error model assumes that ${\bf y}(t)$ is directly observable. It is different from an “output error” model where the recursion has no error, but the recursive signal ${\bf z}(t)$ is hidden and only observed with additive noise (see Fig. 11):
For the output error model, the square error is not a quadratic function of the parameters ${\bf A}$ and ${\bf B}$. Therefore, there is no closed-form solution to the system identification problem, as we had for the equation error model. A few different iterative optimization approaches have been proposed, such as an expectation maximization (EM) algorithm [32], gradient back-propagation through time [33, 34], or “pseudo regression” [1]. The pros and cons of the equation-error versus the output-error models are elaborated in the Discussion section.
Figure 11: Output error model: Here the dynamical variables ${\bf y}(t)$ are not observable. The gray box represents the overall system response $\mathcal{H}$.
Continuing with the examples of the previous section (Fig. 10), despite using the VARX model, when the data was generated with an output error model, the false discovery remains limited at the target of 0.05. However, a common unobserved cause can generate spurious effects $y_{1}\leftarrow y_{2}$. It has been suggested that running a Granger causality model on time-reversed data provides a control for this situation [35, 36]. We have found that running the model on time-reversed data results in spurious effects in all conditions tested here, so it is not clear to us how this can provide a remedy, and the caveat of an unobserved common cause remains when we are not directly observing endogenous variables, but only a noisy version of the internal dynamic.
7.6 Derivation of the De-biased Deviance
To our knowledge, the correction term for the deviance in the case of L2 regularization of a linear estimator is not yet available in the literature. We therefore detail he its derivation.
@@ 7.6 Derivation of the De-biased Deviance
Under the null hypothesis, the de-biased deviance difference can be shown to be asymptotically, as $T\rightarrow\infty$, chi-square distributed with $n$ degrees of freedom, where $n$ is the number of parameters removed in the reduced model and $T$ is the number of observations (i.e., column dimension of $\mathbf{Y}$) [12]. This asymptotic property requires that ridge estimator is consistent. To guarantee this, $\gamma$ must grow slower than $\mathcal{O}(T)$. In practice, fixing $\gamma$ using cross-validation typically meets this condition.